Universal Electronics, Inc. (UEI), which started operations one year ago, has two divisions: Consumer and Commercial. Both divisions invest heavily in R&D, which is assumed to benefit five years. R&D spending is made uniformly throughout the year. UEI has a cost of capital of 11 percent. Selected financial information for the two divisions (in thousands of dollars) for the year just completed follows.

Consumer Commercial
Sales revenue $ 34,000 $ 55,000
Divisional income 6,300 6,450
Divisional investment 30,500 32,250
Current liabilities 2,200 2,000
R&D 2,200 2,200

Required:

Evaluate the performance of the two divisions assuming UEI uses economic value added (EVA).
Answer #1

Ans. : Economic value added:

* It is an estimate of firms economic profit or the value created in excess of required return of company shareholders.

* Economic value added = Net operating profit after tax - (Capital invested * weighed cost of capital)

* Where capital invested = Equity + long term debt, Net operating profit after tax = Earning before interest and tax * ( 1 - Tax rate).

* In EVA, Research and development cost are capitalised which means that these cost are considered as long term assets.

* EVA in case of Consumer division = division income - divisional investment * Cost of capital

= 6300 - 30,500 * 11%

= 6300 - 3355 = 2,945 ( in thousands of dollars).

* EVA in case of commercial decision = 6450 - 32250* 11%

= 6450 - 3547.5

= 2902.5 (in thousands of dollars)

* Performance of consumer division is good as it gives incremental EVA of 42.5 (in thousands of dollars).